OpenStudy (anonymous):

@maggie18854 need help on 2 more??

OpenStudy (anonymous):

ok -2x + 5y = -2 x-3y = 3 what would my answer be but here are my choices... (1,-9) (-9, -4)\ (-9, 0) (0,-4)

OpenStudy (anonymous):

the soultion?

OpenStudy (anonymous):

yes what would be the solution

OpenStudy (anonymous):

kk gimme a min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Next

OpenStudy (anonymous):

x+2y =10 x +y= 6 whts the solution here are my choices again (0,-4) (2,2) (2,4) (-4,-2)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

anything else :)

OpenStudy (anonymous):

NO THANK YOU VERY VERY VERY MUCH LOL :)

OpenStudy (anonymous):

did u get a 100

OpenStudy (anonymous):

np have a nice day :P

OpenStudy (anonymous):

YOU TO :)